OpenStudy (anonymous):

The function H described by H(x)= 2.75+71.48 can be used to predict the height in centimeters, of a women whose humerus (the bone from the elbow to the shoulder) is x cm long. Predict the height of a women whose humerus is 39 cm long . The predicted height of a women whose humerus is 39 cm long is ____ cm

OpenStudy (anonymous):

where is x, in the equation?

OpenStudy (anonymous):

H(x)=2.75x+71.48? or... H(x)=2.75+71.48x? or....?

OpenStudy (anonymous):

H(x)=2.75x+71.48

OpenStudy (anonymous):

We have the equation and we are given x=39. Just plug in x=39 and you get:\[H(39)=2.75(39)+71.48=178.73\]The woman's predicted height is 178.73cm.